One number is 10 more than another. If the sum of twice the smaller number and 3 times the larger number is 55, find the two numbers.

Answer by queenofit(120)

let x = the first number let x+10 = second number
2(x)+3(x+10)= 55
2x+3x+30=55
5x=55-30
5x=25
x=5
x+10=15
check your answer
2(5)+3(5+10)=55
10+15+30=55
55=55